A government agency is investigating a complaint from some concerned citizens who said that there is short-weight selling of rice in a certain town. An agent manufacturer took a random sample of 20 sacks of "50 kilo" sacks of rice from a large shipment and finda that the mean weight is 49.7 kilogram with the standard deviation of 0.35 kilogram. The level of significance is 0.01

1
Expert's answer
2022-05-24T10:08:39-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge50"

"H_a:\mu<50"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=19" degrees of freedom, and the critical value for a left-tailed test is "t_c =-2.539483."The rejection region for this left-tailed test is "R = \{t:t<-2.539483\}."

The t-statistic is computed as follows:



"t=\dfrac{49.7-50}{0.35/\sqrt{20}}=-3.83326"

Since it is observed that "t =-3.83326<-2.539483=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for left-tailed "df=19" degrees of freedom, "t=-3.83326" is "p= 0.000561," and since "p=0.000561<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is less than 50, at the "\alpha = 0.01" significance level.